4.1.
Using one of the laws of exponents, prove that any number raised to the 0'th power equals 1.
thank you.

Answer by stanbon(57967)

Using one of the laws of exponents, prove that any number raised to the 0'th power equals 1.
------------
a/a = 1 as long as a is not zero.
----
But a/a = a^(1-1) = a^0
So a^0 = 1 as long as a is not zero
=======================================
Cheers,
Stan H.